# Mejdl Safran and Imran Ashraf

### v2

@@ Abstract
Method: The method involves the development and implementation of EmoWrite, utilizing a user-centric Recurrent Neural Network (RNN) for thought-to-text conversion. The system incorporates visual feedback and introduces a dynamic keyboard with a contextually adaptive character appearance. Comprehensive evaluation and comparison against existing approaches are conducted, considering various metrics such as accuracy, typing speed, sentiment analysis, emotional state capturing, and user interface latency.
Results: EmoWrite achieves notable results, including a typing speed of 6.6 Words Per Minute (WPM) and 31.9 Characters Per Minute (CPM) with a high accuracy rate of 90.36%. It excels in capturing emotional states, with an Information Transfer Rate (ITR) of 87.55 bits/min for commands and 72.52 bits/min for letters, surpassing other systems. Additionally, it offers an intuitive user interface with low latency of 2.685 seconds.
Conclusion: The introduction of EmoWrite represents a significant stride towards enhancing BCI usability and emotional integration. The findings suggest that EmoWrite holds promising potential for revolutionizing communication aids for individuals with motor disabilities.
@@ 1 Introduction
Mind-reading systems were fiction, but with the advancement of technology, are becoming a reality and helping physically challenged people in performing their tasks e.g., controlling a wheelchair, robotic arm, and cursor. People having a physical disability and speech obstruction are not disabled, rather they are differently abled because they might be impaired with one or two abilities, but their other capabilities can be more precise and accurate as compared to a healthy person 1 . Particularly, when talking about paralytic patients, they might not have an appropriate communication medium to convey their feelings, but their mental activity is more precise if it is utilized efficiently. A lot of research work has been done which is not limited only to assisting in rehabilitation, but to making them self-reliant 2 . Scientists are utilizing brain signals, usually through electroencephalogram (EEG), by extracting useful information from acquired signals. These brain signals are being used in many domains of daily life and exclusively in the medical domain for monitoring alertness, coma or death, brain damage, controlling anesthesia depth, brain development, testing of drugs, and monitoring sleep disorders 2 . EEG signals are also used to resolve speech impediments and eradicate communication barriers of paralytic patients by converting their thoughts (silent speech) to text.
@@
There are two methods used in literature to decode brain signals. The first method directly decodes brain signals into a word, while the second method requires the use of an intermediate output device for converting thought to text. Converting a word directly from the brain to text seems not so much feasible because only limited numbers of words can be interpreted at a time due to the need for additional training, computation power, and resources. Limited information is available regarding the aspect that whether the brain generates the same signals while perceiving similar words or not. Hence, this research area has not yet matured. 3 in their research decode five words only i.e. “Alpha”, “Bravo”, “Charlie”, “Delta” and “Echo” while 4 only decode five characters i.e. a, e, i, o, t. The second method needs a medium, which includes an interface containing characters or words, that can be selected with the help of brain signals. The character selection can be based on two mechanisms i.e. using a virtual keyboard or Visual Evoked Potential (VEP)/Steady-State Visual Evoked Potential (SSVEP) 5 . The virtual keyboard uses raw data or built-in functions of the headset i.e. left, right, up, down, or motor imagery (i.e. imaging movements of hands or feet) 2 , whereas attention-based systems (VEP/SSVEP) focus on some flickering stimulus for selection of characters 5 . The factors which can affect the performance of these systems are related to speed, accuracy, and usability of the system. Moreover, people using these systems are not able to express their feelings accurately because it is problematic to find a proper word to write according to one’s mood, so incorporation of the emotional state along with other commands from the brain will help in better utilization of these systems.
@@
  • Offers BCI-driven remedies for aiding individuals with severe disabilities caused by paralysis and speech impediments.
@@
2 - - - - - - - 95.53% 6.67 CPM
8 - - - - - - 74.95% N/A
16 - - - - - - - - 87.50% N/A
17 - - - - - - - - - N/A 12 WPM
5 - - - - - - - N/A 5-10 CPM(P300) 7.34 CPM (SSVEP) 5 CPM (motor)
18 - - - - - - - - 80-90% 6.5 WPM
19 - - - - - - - - 96% 9 WPM
20 - - - - - - - - - 94% N/A
21 - - - - - - - - N/A 11.93 CPM
22 - - - - - - - - - N/A 9.3 CPM
23 - - - - - - - - - 25% N/A
6 - - - - - - - - N/A N/A
7 - - - - - - - - - N/A N/A
9 - - - - - - - - - N/A 7 CPM
4 - - - - - - - - - - 31% N/A
24 - - - - - - - - - N/A 12 CPM
10 - - - - - - - - - N/A 36 CPM (QWERTY) 39 CPM (OPTI-II) 13.5 CPM (alphabetic)
25 - - - - - - - - - N/A N/A
11 - - - - - - - - - N/A N/A
12 - - - - - - - - - N/A 6.61 CPM
13 - - - - - - - - - - 59.20% N/A
14 - - - - - - - - - N/A N/A
26 - - - - - - - - - - N/A N/A
15 - - - - - - - - - - 99% 5.11 WPM
27 - - - - - - - - - - 79.90% N/A
@@
2 Related Work
To enable communication for paralytic patients some work has been done in the past and still research is ongoing in the domain of BCI-based thoughts-to-text conversion. One of the methods to convert thoughts into text is using a graphical user interface (GUI) consisting of numbers, alphabet, or special characters, which are displayed in a certain order on a virtual keyboard. The brain signals are then used to control the selection of any desired character or alphabet from this virtual keyboard. The major selection methods used in previous systems can be categorized as 1) attention-based control like Visual Evoked Potential (VEP) or Steady-State Visual Evoked Potential (SSVEP), and 2) raw data or built-in functions of the headset to control the cursor or targeted area on the screen. The virtual keyboards are divided into single or multiple layers with static or dynamic keys and their design has a direct influence on the performance of the system. A wide literature survey has been conducted to specify different types of action selection methodologies; character arrangement and virtual keyboard designs.
Some of the major challenges for decoding EEG signals are low signal-to-noise ratio, time consumption, and accuracy. To overcome these challenges, a novel hybrid deep learning approach based on Recurrent Neural Networks (RNNs) and Convolutional Neural Networks (CNNs) is used that converts thoughts to text 2 . EEG signals are used to control the cursor of a personal computer 6 . Features from EEG signals are extracted with a 64-channel NeuroSky headset using a discrete wavelet transform and are classified using machine learning algorithms such as Support Vector Machine (SVM) and Neural Networks (NNs). 7 describe a two-layer hierarchical layout of the keyboard that works with motor imagery signals (left-hand raised, right-hand raised, nodding up ad nodding down).
8 use EEG signals to enhance a written sentence with detected emotions by inserting words in it. Long Short-Term Memory (LSTM) Networks-based language modeling framework is used to verify the sentence correctness by ranking the generated suggestions. ‘Hex-o-spell’, a tilt-based gestural text entry system is introduced by 9 where the letters are arranged in 6 hexagonal shaped boxes that are rearranged after every transition to save time. 4 show how English alphabets are decoded using EEG phase information by using the 64-channel actiCHamp Brain Product to acquire EEG signals. Five alphabets (a, e, l, o, t) are chosen, and against each alphabet, the brain signals are recorded. Results showed that accuracy increased to 31% and the time improved to 200ms. Results prove that most decrypted data lies within the period of 100 to 600ms. High-performance intracortical BCI for communication is described by 10 which provides point-and-click control of the computer. These controls are translated by ReFIT Kalman Filter (RKF) which translates the 2D cursor movement and Hidden Markov Model (HMM) which translates the selection.
@@ 2 Related Work
Figure 1: Basic Flow of Proposed System
3 EmoWrite System
Communication medium plays an important role in Human to Human (H2H) or Human to Machine (H2M) interaction. BCI aids paralytic patients, who cannot communicate, by providing solutions for H2H and H2M interaction. Existing BCI-based work in this domain, especially thought-to-text conversion, is limited in efficiency, accuracy, and number of words per minute. Till now the maximum of 12 WPM has been achieved with a non-invasive technique 17 .
Considering all the discussed challenges, EmoWrite integrates a dynamic keyboard with a circular arrangement of keys. The traversal in the proposed keyboard is controlled by mapping brain commands with facial expressions and using the built-in functions of the headset. It also predicts the next helping verb which is displayed on the right side of the screen. Moreover, the next word prediction is emotion-based as well as personalized. Emotion-based predictions of words assist paralytic patients by efficiently converting their thoughts to text. Furthermore, the machine learning algorithm keeps on retraining itself after a specific interval to predict only the latest and up-to-date words. Additionally, integrating the emotional states of patients with machine learning techniques enhances the performance and productivity of the system. Implementation of EmoWrite aims at reducing the typing delay, increasing accuracy, typing speed, and convenience of the interface. The signals from the brain are acquired through EEG and decoded to convert thoughts to text after extracting information by processing the signals. The extracted information from the brain signals is then classified and mapped to the mental commands (e.g., thinking of left or right direction) or facial expressions (e.g., eye blink, frown, etc.) to perform specific tasks. Emotion state detection has been integrated with machine learning for better productivity of the system. The personalized dynamic arrangement of characters on the screen uses a language model (character sequence pair) and a machine-learning algorithm to show only desired characters on the screen. Finally, the user gets visual feedback through the typed text shown on the screen and the machine learning algorithm also gets feedback from the user to help update its weights for future predictions. EmoWrite is comprised of the following four modules: 1) Data Acquisition 2) Data Processing 3) Basic Cognition, and 4) Communication Interface. The basic flow of the proposed system is shown in Figure 1.
@@ 3.2.1 Emotion Detection
The proposed EmoWrite system utilizes the cortex API’s performance metrics to detect emotions  [^4] https://emotiv.gitbook.io/emotivpro/data_streams/performance-metrics that furnishes crucial insights into a user’s cognitive state through the classification of six key metrics; stress, engagement, interest, excitement, focus, and relaxation. These metrics collectively contribute to the identification of four distinct emotional classes namely, happiness, sadness, anger, and calm. Upon ascertaining the detected emotional class, a correlation finder module identifies a collection of emotion-related words from multiple datasets that have been annotated with emotions 8 , 28 . To ensure data cleanliness, the dataset undergoes preprocessing, which involves the removal of extraneous spaces and symbols. Next, separate Recurrent Neural Network (RNN) models are trained for each emotional class. These models are designed to facilitate emotion-based predictions. The selection of contextualized words depends on the emotional states inferred, and these chosen words are then integrated into the list of predictive words. This comprehensive methodology culminates in the creation of an emotion-sensitive word prediction system, enhancing the accuracy and contextuality of the predicted words.
@@ 3.2.2 User-centric Machine Learning Algorithm
A machine learning algorithm is used along with emotions to predict the next word. EmoWrite employs Recurrent Neural Network (RNN) for predicting contextualized words. RNN has been proven to be the most efficient machine-learning algorithm 29 that provides consistent refinement in the system by requiring less feature engineering, which is a time taking task. It also effectively adopts new data and has parallel processing abilities. RNN has an advantage over other neural networks as all the inputs are dependent on each other, it keeps track of relations with previous words and helps in anticipating preferable output. Being an online algorithm, it updates itself after a specific period when new data is entered by the user.

### v3

# Imran Ashraf

@@ Abstract
Method: The method involves the development and implementation of EmoWrite, utilizing a user-centric Recurrent Neural Network (RNN) for thought-to-text conversion. The system incorporates visual feedback and introduces a dynamic keyboard with a contextually adaptive character appearance. Comprehensive evaluation and comparison against existing approaches are conducted, considering various metrics such as accuracy, typing speed, sentiment analysis, emotional state capturing, and user interface latency. The data required for this experiment was obtained from a total of 72 volunteers (40 male and 32 female) aged between 18 and 40.
Results: EmoWrite achieves notable results, including a typing speed of 6.6 Words Per Minute (WPM) and 31.9 Characters Per Minute (CPM) with a high accuracy rate of 90.36%. It excels in capturing emotional states, surpassing other systems with an Information Transfer Rate (ITR) of 87.55 bits/min for commands and 72.52 bits/min for letters. It also offers an intuitive user interface with a low latency of 2.685 seconds.
Conclusion: The introduction of EmoWrite represents a significant stride towards enhancing BCI usability and emotional integration. The findings indicate that EmoWrite shows promising potential in enhancing communication methods, with future implications for individuals with motor disabilities.
@@ 1 Introduction
Mind-reading systems were fiction, but with the advancement of technology, are becoming a reality and helping physically challenged people in performing their tasks e.g., controlling a wheelchair, robotic arm, and cursor. People having a physical disability and speech obstruction are not disabled, rather they are differently abled because they might be impaired with one or two abilities, but their other capabilities can be more precise and accurate as compared to a healthy person 1 . Particularly, when talking about paralytic patients, might not have an appropriate communication medium to convey their feelings, but their mental activity is more precise if it is utilized efficiently. A lot of research work has been done which is not limited only to assisting in rehabilitation, but to making them self-reliant 2 . Scientists are utilizing brain signals, usually through electroencephalogram (EEG), by extracting useful information from acquired signals. These brain signals are being used in many domains of daily life and exclusively in the medical domain for monitoring alertness, coma or death, brain damage, controlling anesthesia depth, brain development, testing of drugs, and monitoring sleep disorders 2 . EEG signals are also used to resolve speech impediments and eradicate communication barriers of paralytic patients by converting their thoughts (silent speech) to text.
@@
There are two methods used in literature to decode brain signals. The first method directly decodes brain signals into a word, while the second method requires the use of an intermediate output device for converting thought to text. Converting a word directly from the brain to text seems not so feasible because only limited numbers of words can be interpreted at a time due to the need for additional training, computation power, and resources. Limited information is available regarding the aspect that whether the brain generates the same signals while perceiving similar words or not. Hence, this research area has not yet matured. 3 in their research decode five words only i.e. “Alpha”, “Bravo”, “Charlie”, “Delta” and “Echo” while 4 only decode five characters i.e. a, e, i, o, t. The second method needs a medium, which includes an interface containing characters or words, that can be selected with the help of brain signals. The character selection can be based on two mechanisms i.e. using a virtual keyboard or Visual Evoked Potential (VEP)/Steady-State Visual Evoked Potential (SSVEP) 5 . The virtual keyboard uses raw data or built-in functions of the headset i.e. left, right, up, down, or motor imagery (i.e. imaging movements of hands or feet) 2 , whereas attention-based systems (VEP/SSVEP) focus on some flickering stimulus for selection of characters 5 . The factors that can affect the performance of these systems are related to speed, accuracy, and usability of the system. Moreover, people using these systems are not able to express their feelings accurately because it is problematic to find a proper word to write according to one’s mood, so incorporation of the emotional state along with other commands from the brain will help in better utilization of these systems.
Using a deep learning approach can be significantly important in this regard and make potential contributions. For example, the study uppal2023enhancing introduced a technique to predict brain strokes with high accuracy. The brain stroke data was used to build the model which is a multilayer perceptron (MLP). It employed multiple optimizers, including adaptive moment estimation with maximum, root mean squared propagation (RMSProp), and the adaptive learning rate method. Experimental results indicate that the MLP model combined with the RMSProp optimizer performed the best, achieving a training accuracy of 95.8% and a testing accuracy of 94.9%. Another possible venture is the self-supervised learning phenomenon where large-sized unlabeled datasets are used for model training. The existing research is lacking on self-supervised learning, the study abdulrazzaq2024consequential discusses applications of self-supervised learning from the prospect of industrial engineering and medicine. The authors identify the key possibilities for prediction in these fields using self-supervised learning. Analysis is carried out in the context of medical staff predicting patient ailments more efficiently, without relying on traditional numerical models that require a lot of computation, time, storage, and effort for data annotation. Similarly, functional near-infrared spectroscopy (fNIRS) has been used for brain-computer interface (BCI) tasks. The study zafar2023metaheuristic introduces a method for selecting important features for brain-computer interface (BCI) applications using functional near-infrared spectroscopy (fNIRS). Temporal statistical features, like the mean, slope, maximum, skewness, and kurtosis, were calculated from all channels to create a training vector. Seven different optimization algorithms were tested for their ability to classify data using a k-nearest neighbor cost function: particle swarm optimization, cuckoo search optimization, the firefly algorithm, the bat algorithm, flower pollination optimization, whale optimization, and grey wolf optimization (GWO). This method was tested on an online dataset of motor imagery (MI) and mental arithmetic (MA) tasks from 29 healthy subjects. The results showed that using the features selected by these optimization algorithms significantly improved classification accuracy compared to using all available features.
@@
  • Introduces BCI-driven solutions with the potential to support individuals with severe disabilities, focusing on initial validation and future applications for those with paralysis and speech impairments.
@@
2 Related Work
To enable communication for paralytic patients some work has been done in the past and still research is ongoing in the domain of BCI-based thoughts-to-text conversion. One of the methods to convert thoughts into text is using a graphical user interface (GUI) consisting of numbers, alphabet, or special characters, which are displayed in a certain order on a virtual keyboard.
@@ 2 Related Work
Zhang et al. (2018) 2 - - - - - - - 95.53% 6.67 CPM
Gupta et al. (2019) 8 - - - - - - 74.95% N/A
Masud et al. (2017) 16 - - - - - - - - 87.50% N/A
Chen et al. (2015) 17 - - - - - - - - - N/A 12 WPM
Cecotti (2011) 5 - - - - - - - N/A 5-10 CPM(P300) 7.34 CPM (SSVEP) 5 CPM (motor)
Chen et al. (2014) 18 - - - - - - - - 80-90% 6.5 WPM
Spüler et al. (2012) 19 - - - - - - - - 96% 9 WPM
Higger et al. (2016) 20 - - - - - - - - - 94% N/A
Akce et al. (2014) 21 - - - - - - - - N/A 11.93 CPM
Cecotti (2016) 22 - - - - - - - - - N/A 9.3 CPM
Ben-Ami et al. (2019) 23 - - - - - - - - - 25% N/A
Alomari et al. (2014) 6 - - - - - - - - N/A N/A
Hayet et al. (2019) 7 - - - - - - - - - N/A N/A
Williamson et al. (2009) 9 - - - - - - - - - N/A 7 CPM
Wang et al. (2018) 4 - - - - - - - - - - 31% N/A
Jarosiewicz et al. (2015) 24 - - - - - - - - - N/A 12 CPM
Pandarinath et al. (2017) 10 - - - - - - - - - N/A 36 CPM (QWERTY) 39 CPM (OPTI-II) 13.5 CPM (alphabetic)
Arijit et al. (2013) 25 - - - - - - - - - N/A N/A
Topal et al. (2012) 11 - - - - - - - - - N/A N/A
Pathirana et al. (2018) 12 - - - - - - - - - N/A 6.61 CPM
Andi et al. (2018) 13 - - - - - - - - - - 59.20% N/A
Birbaumer et al. (2000) 14 - - - - - - - - - N/A N/A
George et al. (2014) 26 - - - - - - - - - - N/A N/A
Mackenzie et al. (2010) 15 - - - - - - - - - - 99% 5.11 WPM
Morooka et al. (2018) 27 - - - - - - - - - - 79.90% N/A
@@
The brain signals are then used to control the selection of any desired character or alphabet from this virtual keyboard. The major selection methods used in previous systems can be categorized as 1) attention-based control like Visual Evoked Potential (VEP) or Steady-State Visual Evoked Potential (SSVEP), and 2) raw data or built-in functions of the headset to control the cursor or targeted area on the screen. The virtual keyboards are divided into single or multiple layers with static or dynamic keys and their design has a direct influence on the performance of the system. A wide literature survey has been conducted to specify different types of action selection methodologies; character arrangement and virtual keyboard designs.
Some of the major challenges for decoding EEG signals are low signal-to-noise ratio, time consumption, and accuracy. To overcome these challenges, a novel hybrid deep learning approach based on Recurrent Neural Networks (RNNs) and Convolutional Neural Networks (CNNs) is used that converts thoughts to text 2 . EEG signals are used to control the cursor of a personal computer 6 . Features from EEG signals are extracted with a 64-channel NeuroSky headset using a discrete wavelet transform and are classified using machine learning algorithms such as Support Vector Machine (SVM) and Neural Networks (NNs). 7 describe a two-layer hierarchical layout of the keyboard that works with motor imagery signals (left-hand raised, right-hand raised, nodding up and nodding down).
8 use EEG signals to enhance a written sentence with detected emotions by inserting words in it. Long Short-Term Memory (LSTM) Networks-based language modeling framework is used to verify the sentence correctness by ranking the generated suggestions. ‘Hex-o-spell’, a tilt-based gestural text entry system is introduced by 9 where the letters are arranged in 6 hexagonal shaped boxes that are rearranged after every transition to save time. 4 show how English alphabets are decoded using EEG phase information by using the 64-channel actiCHamp Brain Product to acquire EEG signals. Five alphabets (a, e, l, o, t) are chosen, and against each alphabet, the brain signals are recorded. Results showed that accuracy increased to 31% and the time improved to 200ms. Results prove that most decrypted data lies within the period of 100 to 600ms. High-performance intracortical BCI for communication is described by 10 which provides point-and-click control of the computer. These controls are translated by the ReFIT Kalman Filter (RKF) which translates the 2D cursor movement and the Hidden Markov Model (HMM) which translates the selection.
@@
3 Methods
Communication medium plays an important role in human-to-human (H2H) or human-to-machine (H2M) interaction. BCI aids paralytic patients, who cannot communicate, by providing solutions for H2H and H2M interaction. Existing BCI-based work in this domain, especially thought-to-text conversion, is limited in efficiency, accuracy, and number of words per minute. Till now the maximum of 12 WPM has been achieved with a non-invasive technique 17 .
Considering all the discussed challenges, EmoWrite integrates a dynamic keyboard with a circular arrangement of keys. The traversal in the proposed keyboard is controlled by mapping brain commands with facial expressions and using the built-in functions of the headset. It also predicts the next helping verb which is displayed on the right side of the screen. Moreover, the next word prediction is emotion-based as well as personalized. Emotion-based predictions of words assist paralytic patients by efficiently converting their thoughts to text. Furthermore, the machine learning algorithm keeps on retraining itself after a specific interval to predict only the latest and up-to-date words. Additionally, integrating the emotional states of patients with machine learning techniques enhances the performance and productivity of the system.
Implementation of EmoWrite aims at reducing the typing delay, increasing accuracy, typing speed, and convenience of the interface. The signals from the brain are acquired through EEG and decoded to convert thoughts to text after extracting information by processing the signals. The extracted information from the brain signals is then classified and mapped to the mental commands (e.g., thinking of left or right direction) or facial expressions (e.g., eye blink, frown, etc.) to perform specific tasks. Emotion state detection has been integrated with machine learning for better productivity of the system. The personalized dynamic arrangement of characters on the screen uses a language model (character sequence pair) and a machine-learning algorithm to show only the desired characters on the screen. Finally, the user gets visual feedback through the typed text shown on the screen and the machine learning algorithm also gets feedback from the user to help update its weights for future predictions. EmoWrite is comprised of the following four modules: 1) Data Acquisition 2) Data Processing 3) Basic Cognition, and 4) Communication Interface. The basic flow of the proposed system is shown in Figure 1.
@@ 3.1 Data Acquisition
The data for this experiment was collected from 72 volunteers (55.6% male and 44.4% female), with a mean age of 29 years and a standard deviation of 6.5 years, at the Advanced Communication Networks Lab at COMSATS University Islamabad, Lahore Campus. None of the participants had any previous BCI experience. Before the data collection process, informed consent was obtained from all participants and the study protocol was approved and supervised by the Ethics Committee of COMSATS University Islamabad, Lahore Campus. All experiments were performed under relevant guidelines and regulations. The duration of the study was two years from 2019 to 2021.
@@ 3.2 Data Processing
(a) First Set of Characters.
(b) Character set after selection of ’t’.
Figure 2: Keyboard Interface.
@@ 3.2.1 Emotion Detection
The proposed EmoWrite system utilizes the cortex API’s performance metrics to detect emotions  [^4] https://emotiv.gitbook.io/emotivpro/data_streams/performance-metrics that furnishes crucial insights into a user’s cognitive state through the classification of six key metrics; stress, engagement, interest, excitement, focus, and relaxation. These metrics collectively contribute to the identification of four distinct emotional classes namely, happiness, sadness, anger, and calm. Upon ascertaining the detected emotional class, a correlation finder module identifies a collection of emotion-related words from multiple datasets that have been annotated with emotions 8 , 28 . To ensure data cleanliness, the dataset undergoes preprocessing, which involves the removal of extraneous spaces and symbols. Next, separate Recurrent Neural Network (RNN) models are trained for each emotional class. These models are designed to facilitate emotion-based predictions. The selection of contextualized words depends on the emotional states inferred, and these chosen words are then integrated into the list of predictive words. This comprehensive methodology leads to an emotion-sensitive word prediction system, enhancing the accuracy and contextuality of the predicted words.
@@ 3.2.2 User-centric Machine Learning Algorithm
A machine learning algorithm is used along with emotions to predict the next word. EmoWrite employs a Recurrent Neural Network (RNN) for predicting contextualized words. RNN has been proven to be the most efficient machine-learning algorithm 29 that provides consistent refinement in the system by requiring less feature engineering, which is a time-consuming task. It also effectively adopts new data and has parallel processing abilities. RNN has an advantage over other neural networks as all the inputs are dependent on each other, it keeps track of relations with previous words and helps in anticipating preferable output. Being an online algorithm, it updates itself after a specific period when new data is entered by the user.